OpenStudy (anonymous):

the equation x2+y2+ax+by+c=0 in standard form

OpenStudy (ranga):

x2+y2+ax+by+c=0 group the c terms and y terms separately and complete the squares: (x^2 + ax) + (y^2 + by) + c = 0 complete the square for the terms inside the parenthesis.

OpenStudy (anonymous):

can you show me?

OpenStudy (ranga):

(x^2 + ax) + (y^2 + by) + c = 0 { (x + a/2)^2 - a^2/4 } + { (y + b/2)^2 - b^2/4 } + c = 0 (x + a/2)^2 - a^2/4 + (y + b/2)^2 - b^2/4 + c = 0 (x+a/2)^2 + (y+b/2)^2 - (a^2 + b^2 - 4c)/4 = 0 (x+a/2)^2 + (y+b/2)^2 = (a^2 + b^2 - 4c)/4 (x+a/2)^2 + (y+b/2)^2 = { sqrt((a^2 + b^2 - 4c))/2 }^2